Please explain how I can find the inverse (f-1)
ex: y=2x+6
Thank you

Answer by nerdybill(7384)

Please explain how I can find the inverse (f-1)
ex: y=2x+6
.
Switch the variables; substitute x for y and y for x:
x=2y+6
Solve for y:
x-6=2y
x/2 - 3 = y
.
Inverse is:
y^(-1) = x/2 - 3